# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ IV.2.1 Second-order dynamical equations: non-commutative formulation
From expression (55), assuming the gauge fixing condition $N=1$, the following dynamical equations for $\eta(\tau)$ and $\xi(\eta)$ result:
$\dot{\eta}=\frac{\partial{\cal H}^{NC}}{\partial p_{\eta}}=-\frac{p_{\eta}}{\eta}-\frac{\gamma}{2\eta^{3\alpha}}\,;$ (56)
$\dot{\xi}=\frac{\partial{\cal H}^{NC}}{\partial p_{\xi}}=\frac{1}{2\eta^{3\alpha}}\,;$ (57)
$\quad\dot{p}_{\xi}=-\frac{\partial{\cal H}^{NC}}{\partial{\xi}}=-\frac{\alpha}{2\eta^{3\alpha}}\,;$ (58)
@@
[…]
@@
$\dot{p}_{\eta}=\frac{\partial p_{\eta}}{\partial t}=-\frac{\partial(\eta\dot{\eta})}{\partial t}=-\dot{\eta}^{2}-\eta\ddot{\eta},$ (61)
Equation (59) combined with (61) may be recast in the form
$2\eta(t)\ddot{\eta}(t)+\dot{\eta}^{2}(t)+\frac{3\alpha\gamma\dot{\eta}}{\eta^{3\alpha}}+V(\eta,t)=0\,,$ (62)
@@
$V(\eta,t)\!=\!g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}-\frac{3\alpha^{2}t}{\eta^{6\alpha-1}}\,.$ (63)
Figures 8 show 3D graphical illustrations of the potential $V(\eta,t)$ of equation (63) and the corresponding contour plots characterizing the dependence of the scale factor $\eta(t)$ on time. The contour plots in the commutative formulation, as seen in figure 6, exhibit a symmetrical distribution of intensities or amplitudes of the potential $V(u,t)$ intensities or amplitudes around the lines $t=0$ and $u(t)=0$. In the non-commutative formulation, this symmetry is broken, indicating a mixture of intensities or amplitudes of the potential $V(\eta,t)$. The potential $V(\eta,t)$ simulates the presence of different compositions of matter in the primordial Universe that imply structural modifications of the spacetime structure, shaping this way its curvature that depends locally on the amount and distribution of matter or, equivalently, energy. This symmetry breaking reveals the potentiality of a non-commutative formulation in terms of its implications in affecting not only the curvature of space-time, but furthermore, the capture of short and long scales, boosting the evolution dynamics of the wave function of the Universe and the cosmic scale factor. Insofar, as the presence of the potential is associated with a force, of a gravitational character, which may constitute the propelling element of the acceleration of the primordial Universe. These results provide conceptual elements that indicate a reconfiguration of matter on small scales of dimensions. This reconfiguration of matter, as previously observed, represents a typical characteristic of a non-commutative formulation, which drives, through a non-symmetric gravitational force, the acceleration of the Universe. Furthermore, due to the dual role of gravity, spacetime is both warped by matter and matter in turn experiences a warped spacetime due to the presence of other distributions of matter, which intensifies the effects of higher curvature terms.
@@
Figure 9: Typical sample family solutions for the scale factor of the branch-cut Universe, $\eta(t)$, using the second-order time derivative non-commutative approach equation (62). The solutions are for varying initial conditions $\eta(0)$ and $\eta^{\prime}(0)$. In the left figure (a), the values of the running coupling constants and parameters are as follows: $\tilde{g}_{r}=-0.4$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=-0.333$; $g_{s}=-0.03$; $\alpha=-1/2$; $\gamma=1$; ${\cal C}=1$. In figure (b), the values of the running coupling constants and parameters remain the same, except for $\alpha=-3/4$; $\gamma=-1$. In figure (c), the values for $\alpha$ and $\gamma$ are $\alpha=-3/4$ and $\gamma=-1/3$.
IV.2.2 First-order dynamical equations: non-commutative formulation
@@
Upon integrating equation (61), while assuming that $\eta$ possesses explicit time dependence allowing for separation from time, and further considering that $\dot{\eta}$ and $\eta$ function as independent variables, we obtain
$\dot{\eta}+v(\eta,t)=0\,,$ (64)
@@
$v(\eta,t)=\frac{1}{3\eta}\Biggl{\{}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{(}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{)}t-\frac{3}{2}\frac{\alpha^{2}t^{2}}{\eta^{6\alpha-1}}\Biggr{\}}\,.$ (65)
Figure 10 presents graphical representations of the potential $v(\eta,t)$ from equation (65), indicating the presence of a torsion deformation. More specifically, it indicates a deformation that resembles a kind of torsional conformation of an object due to the application of an “external” torque.
IV.2.3 Torsion deformation
A widely used method for constructing non-commutative spacetime algebraic structures and non-commutative field theories involves the concept of torsion (or twist) deformation. The graphical representations of the potential $v(\eta,t)$ from equation (65) and the accompanying contour plots in figure 10 in turn resemble, in particular, a characteristic torsion (or twist) deformation of the spacetime geometry. In the following, we focus on the perspectives and implications of this type of space-time distortion within the context of a background branch cut.
Figure 10: Graphical illustrations of the potential $v(\eta,t)$ of equation (65) and contour plots. The values of the running coupling constants and the parameters of the non-commutative algebra for the figures are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\alpha=-5/6$; $\gamma=3/5$ ; ${\cal C}=1$.
@@ IV.2.3 Torsion deformation
$\Gamma^{\alpha}_{\mu\nu}=\tilde{\Gamma}^{\alpha}_{\mu\nu}+\Omega^{\alpha}_{\mu\nu}$ (66)
@@
$\Omega^{\alpha}_{\mu\nu}=S^{\alpha}_{\mu\nu}+S_{\mu\nu}^{\alpha}+S_{\nu\mu}^{\alpha}$ (67)
@@
${\cal F}=e^{-\frac{i}{2}\theta^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}}\,;\quad{\cal F}^{-1}=e^{\frac{i}{2}\theta^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}}\,.$ (68)
@@
$f\star g:=\mu\circ{\cal F}^{-1}(f\otimes g)\,$ (69)
@@
${\cal F}=f^{\alpha}\otimes f_{\alpha}\,;\quad\mbox{with}\quad{\cal F}^{-1}=\bar{f}^{\alpha}\otimes\bar{f}_{\alpha}\,,$ (70)
@@
$f\star g:=\bar{f}^{\alpha}(f)\bar{f}_{\alpha}(g)\,.$ (71)
@@
$(\mathbf{x},\mathbf{y})\mapsto\mu_{\star}(\mathbf{x},\mathbf{y})(x,y)=\mu(\bar{f}^{\alpha}(\mathbf{x},\bar{f}^{\alpha}(\mathbf{x}).$ (73)
@@
IV.2.4 Results: non-commutative dynamical equations
Figure 9 presents typical sample family solutions of the scale factor of the branch-cut Universe, $\eta(t)$, obtained through the combination of second-order in time $\ddot{\eta}(t)$ and first-order in $\dot{\eta}(t)$ non-commutative approach given by equation (62). These solutions are for variations in the initial conditions $\eta(0)$ and $\eta^{\prime}(0)$. Figure 11 shows typical sample family solutions of $\eta(t)$ using the first-order in time $\dot{\eta}$ non-commutative approach given by equation (64), with variations in the value of $\eta(0)$. These results demonstrate the significant impact of the non-commutative formulation on the dynamics of the early Universe’s evolution. The findings reveal that the behavior of the scale factor $\eta(t)$ is notably different from what is predicted in a commutative formulation, particularly concerning the acceleration components of the early Universe. While the Universe may experience a deceleration phase in its initial moments of expansion for specific parameter sets, it abruptly undergoes a drastic acceleration. This acceleration is characterized by an evolutionary curve that approaches a ninety-degree angle with the x-axis, nearly parallel to the y-axis. These results align with one of the key propositions of this work, which aims to understand the mechanisms driving the acceleration of the Universe’s expansion.
@@ IV.2.4 Results: non-commutative dynamical equations
V Gravitational Waves from the Primordial Universe
@@ V.1 Linearized BCG equations
$g^{[\rm ac]}={\zeta}^{[\rm ac]}_{\mu\nu}+{h}^{[\rm ac]}_{\mu\nu}\,,\quad\mbox{with}\quad|{h}^{[\rm ac]}_{\mu\nu}|<<1\,.$ (74)
@@
Similarly to General Relativity, the branch-cut gravity equations relate the spacetime geometry, encoded in the metric $g^{[\rm ac]}$, to matter described by the energy-momentum tensor $T^{[\rm ac]}_{\mu\nu}$ analytically continued to the complex plane. The Ricci tensor $R^{[\rm ac]}_{\mu\nu}$ and Ricci scalar $R^{[\rm ac]}$ for the linearized theory are computed following the usual scheme, starting from the analytically continued Christoffel symbol which will lead to an analytically continued Riemann curvature tensor. From equation (74), following similar mathematical procedures of the standard theory, we obtain for the analytically continued linearized Christoffel symbol
$\Gamma^{{[\rm ac]}\rho}_{\mu\nu}$ $=$ $\frac{1}{2}g^{{[\rm ac]}\rho\sigma}\Bigl{[}\partial_{\mu}g^{[\rm ac]}_{\nu\sigma}+\partial_{\nu}g^{[\rm ac]}_{\mu\sigma}-\partial_{\sigma}g^{[\rm ac]}_{\mu\nu}\Bigr{]}$ (76)
@@
$R^{{[\rm ac]}\mu}_{\nu\sigma\rho}$ $=$ $\partial_{\sigma}\Gamma^{{[\rm ac]}\mu}_{\nu\rho}-\partial_{\rho}\Gamma^{{[\rm ac]}\mu}_{\nu\sigma}+\Gamma^{{[\rm ac]}\mu}_{\sigma\lambda}\Gamma^{{[\rm ac]}\lambda}_{\nu\rho}+\Gamma^{{[\rm ac]}\mu}_{\rho\lambda}\Gamma^{{[\rm ac]}\lambda}_{\nu\sigma}\,,$ (77)
@@
$R^{[\rm ac]}_{\mu\nu}$ $=$ $R^{{[\rm ac]}\rho}_{\mu\rho\nu}$ (78)
$=$ $\frac{1}{2}\Bigl{[}\partial_{\rho}\partial_{\mu}h^{{[\rm ac]}\rho}_{\nu}+\partial_{\rho}\partial_{\nu}h^{{[\rm ac]}\rho}_{\mu}-\partial_{\mu}\partial_{\nu}h^{[\rm ac]}-\Box h_{\mu\nu}\Bigr{]}$
$+$ ${\cal O}({h^{[\rm ac]}}^{2})\,.$
Similarly, the the Ricci scalar is
$R^{[\rm ac]}=g^{{[\rm ac]}\mu\nu}R^{[\rm ac]}_{\mu\nu}=\partial_{\mu}\partial_{\nu}h^{{[\rm ac]}\mu\nu}-\Box h^{[\rm ac]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (79)
@@
$G^{[\rm ac]}_{\mu\nu}=-\frac{1}{2}\Bigl{[}\Box h^{[\rm ac]}_{\mu\nu}+\zeta^{[\rm ac]}_{\mu\nu}\partial^{\rho}\partial^{\sigma}h^{[\rm ac]}_{\rho\sigma}-\zeta^{[\rm ac]}_{\mu\nu}\Box h^{[\rm ac]}-\partial^{\rho}\partial_{\nu}h^{[\rm ac]}_{\mu\rho}-\partial_{\rho}\partial_{\mu}h^{{[\rm ac]}\rho}_{\nu}+\partial_{\nu}\partial_{\mu}h\Bigr{]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (80)
@@
$G^{[\rm ac]}_{\mu\nu}=-\frac{1}{2}\Bigl{[}\Box\bar{h}^{[\rm ac]}_{\mu\nu}+\zeta^{[\rm ac]}_{\mu\nu}\partial^{\rho}\partial^{\sigma}\bar{h}^{[\rm ac]}_{\rho\sigma}-\partial^{\rho}\partial_{\nu}\bar{h}^{[\rm ac]}_{\mu\rho}-\partial^{\rho}\partial_{\mu}\bar{h}^{{[\rm ac]}}_{\nu\rho}\Bigr{]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (81)
V.2 Stochastic gravitational wave background

### 03 Feb 2024 v2

@@ 4.2.1 Second-order dynamical equations: non-commutative formulation
From expression (4.16), assuming the gauge fixing condition $N=1$, the following dynamical equations for $\eta(\tau)$ and $\xi(\eta)$ result:
$\dot{\eta}$ $=$ $\frac{\partial{\cal H}^{NC}}{\partial p_{\eta}}=-\frac{p_{\eta}}{\eta}-\frac{\gamma}{2\eta^{3\alpha}}\,;$ (4.17)
$\dot{\xi}$ $=$ $\frac{\partial{\cal H}^{NC}}{\partial p_{\xi}}=\frac{1}{2\eta^{3\alpha}}\,;$ (4.18)
$\dot{p}_{\xi}$ $=$ $-\frac{\partial{\cal H}^{NC}}{\partial{\xi}}=-\frac{\alpha}{2\eta^{3\alpha}}\,;$ (4.19)
@@
$=$ $\!-\frac{1}{2\eta}\Biggl{[}\frac{1}{\eta}p^{2}_{\eta}+\frac{3\alpha\gamma}{\eta^{3\alpha}}p_{\eta}-\frac{g_{r}}{\eta}-g_{k}\eta-2g_{q}\eta^{2}\!-\!3g_{\Lambda}\eta^{3}-3\frac{g_{s}}{\eta^{3}}-\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha-1}}+\frac{3\alpha^{2}t}{\eta^{6\alpha-1}}\Biggr{]}\,,$
@@
$\dot{p}_{\eta}=\frac{\partial p_{\eta}}{\partial t}=-\frac{\partial(\eta\dot{\eta})}{\partial t}=-\dot{\eta}^{2}-\eta\ddot{\eta},$ (4.22)
Equation (4.20) combined with (4.22) may be recast in the form
$2\eta(t)\ddot{\eta}(t)+\dot{\eta}^{2}(t)+\frac{3\alpha\gamma\dot{\eta}}{\eta^{3\alpha}}+V(\eta,t)=0\,,$ (4.23)
@@
$V(\eta,t)\!=\!g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}-\frac{3\alpha^{2}t}{\eta^{6\alpha-1}}\,.$ (4.24)
Figures 8 show 3D graphical illustrations of the potential $V(\eta,t)$ of equation (4.24) and the corresponding contour plots characterizing the dependence of the scale factor $\eta(t)$ on time. The contour plots in the commutative formulation, as seen in figure 6, exhibit a symmetrical distribution of intensities or amplitudes of the potential $V(u,t)$ intensities or amplitudes around the lines $t=0$ and $u(t)=0$. In the non-commutative formulation, this symmetry is broken, indicating a mixture of intensities or amplitudes of the potential $V(\eta,t)$. The potential $V(\eta,t)$ simulates the presence of different compositions of matter in the primordial Universe that imply structural modifications of the spacetime structure, shaping this way its curvature that depends locally on the amount and distribution of matter or, equivalently, energy. This symmetry breaking reveals the potentiality of a non-commutative formulation in terms of its implications in affecting not only the curvature of space-time, but furthermore, the capture of short and long scales, boosting the evolution dynamics of the wave function of the Universe and the cosmic scale factor. Insofar, as the presence of the potential is associated with a force, of a gravitational character, which may constitute the propelling element of the acceleration of the primordial Universe. These results provide conceptual elements that indicate a reconfiguration of matter on small scales of dimensions. This reconfiguration of matter, as previously observed, represents a typical characteristic of a non-commutative formulation, which drives, through a non-symmetric gravitational force, the acceleration of the Universe. Furthermore, due to the dual role of gravity, spacetime is both warped by matter and matter in turn experiences a warped spacetime due to the presence of other distributions of matter, which intensifies the effects of higher curvature terms.
@@
Figure 9: Typical sample family solutions for the scale factor of the branch-cut Universe, $\eta(t)$, using the second-order time derivative non-commutative approach equation (4.23). The solutions are for varying initial conditions $\eta(0)$ and $\eta^{\prime}(0)$. In the left figure (a), the values of the running coupling constants and parameters are as follows: $\tilde{g}_{r}=-0.4$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=-0.333$; $g_{s}=-0.03$; $\alpha=-1/2$; $\gamma=1$; ${\cal C}=1$. In figure (b), the values of the running coupling constants and parameters remain the same, except for $\alpha=-3/4$; $\gamma=-1$. In figure (c), the values for $\alpha$ and $\gamma$ are $\alpha=-3/4$ and $\gamma=-1/3$.
4.2.2 First-order dynamical equations: non-commutative formulation
@@
Upon integrating equation (4.22), while assuming that $\eta$ possesses explicit time dependence allowing for separation from time, and further considering that $\dot{\eta}$ and $\eta$ function as independent variables, we obtain
$\dot{\eta}+v(\eta,t)=0\,,$ (4.25)
@@
$v(\eta,t)=\frac{1}{3\eta}\Biggl{\{}-\frac{3\alpha\gamma}{\bigl{(}3\alpha-1\bigr{)}\eta^{3\alpha-1}}+\Biggl{(}g_{k}+2g_{q}\eta-3g_{\Lambda}\eta^{2}+\frac{g_{r}}{\eta^{2}}+3\frac{g_{s}}{\eta^{4}}+\frac{\alpha\bigl{(}3\alpha-1\bigr{)}}{\eta^{3\alpha}}\Biggr{)}t-\frac{3}{2}\frac{\alpha^{2}t^{2}}{\eta^{6\alpha-1}}\Biggr{\}}\,.$ (4.26)
Figure 10 presents graphical representations of the potential $v(\eta,t)$ from equation (4.26), indicating the presence of a torsion deformation. More specifically, it indicates a deformation that resembles a kind of torsional conformation of an object due to the application of an “external” torque.
4.2.3 Torsion deformation
A widely used method for constructing non-commutative spacetime algebraic structures and non-commutative field theories involves the concept of torsion (or twist) deformation. The graphical representations of the potential $v(\eta,t)$ from equation (4.26) and the accompanying contour plots in figure 10 in turn resemble, in particular, a characteristic torsion (or twist) deformation of the spacetime geometry. In the following, we focus on the perspectives and implications of this type of space-time distortion within the context of a background branch cut.
Figure 10: Graphical illustrations of the potential $v(\eta,t)$ of equation (4.26) and contour plots. The values of the running coupling constants and the parameters of the non-commutative algebra for the figures are: $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{r}=0.4$; $g_{s}=-0.03$; $\alpha=-5/6$; $\gamma=3/5$ ; ${\cal C}=1$.
@@ 4.2.3 Torsion deformation
$\Gamma^{\alpha}_{\mu\nu}=\tilde{\Gamma}^{\alpha}_{\mu\nu}+\Omega^{\alpha}_{\mu\nu}$ (4.27)
@@
$\Omega^{\alpha}_{\mu\nu}=S^{\alpha}_{\mu\nu}+S_{\mu\nu}^{\alpha}+S_{\nu\mu}^{\alpha}$ (4.28)
@@
${\cal F}=e^{-\frac{i}{2}\theta^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}}\,;\quad{\cal F}^{-1}=e^{\frac{i}{2}\theta^{\mu\nu}\partial_{\mu}\otimes\partial_{\nu}}\,.$ (4.29)
@@
$f\star g:=\mu\circ{\cal F}^{-1}(f\otimes g)\,$ (4.30)
@@
${\cal F}=f^{\alpha}\otimes f_{\alpha}\,;\quad\mbox{with}\quad{\cal F}^{-1}=\bar{f}^{\alpha}\otimes\bar{f}_{\alpha}\,,$ (4.31)
@@
$f\star g:=\bar{f}^{\alpha}(f)\bar{f}_{\alpha}(g)\,.$ (4.32)
@@
$(\mathbf{x},\mathbf{y})\mapsto\mu_{\star}(\mathbf{x},\mathbf{y})(x,y)=\mu(\bar{f}^{\alpha}(\mathbf{x},\bar{f}^{\alpha}(\mathbf{x}).$ (4.34)
@@
4.2.4 Results: non-commutative dynamical equations
Figure 9 presents typical sample family solutions of the scale factor of the branch-cut Universe, $\eta(t)$, obtained through the combination of second-order in time $\ddot{\eta}(t)$ and first-order in $\dot{\eta}(t)$ non-commutative approach given by equation (4.23). These solutions are for variations in the initial conditions $\eta(0)$ and $\eta^{\prime}(0)$. Figure 11 shows typical sample family solutions of $\eta(t)$ using the first-order in time $\dot{\eta}$ non-commutative approach given by equation (4.25), with variations in the value of $\eta(0)$. These results demonstrate the significant impact of the non-commutative formulation on the dynamics of the early Universe’s evolution. The findings reveal that the behavior of the scale factor $\eta(t)$ is notably different from what is predicted in a commutative formulation, particularly concerning the acceleration components of the early Universe. While the Universe may experience a deceleration phase in its initial moments of expansion for specific parameter sets, it abruptly undergoes a drastic acceleration. This acceleration is characterized by an evolutionary curve that approaches a ninety-degree angle with the x-axis, nearly parallel to the y-axis. These results align with one of the key propositions of this work, which aims to understand the mechanisms driving the acceleration of the Universe’s expansion.
@@ 4.2.4 Results: non-commutative dynamical equations
5 Gravitational Waves from the Primordial Universe
@@ 5.1 Linearized BCG equations
$g^{[\rm ac]}={\zeta}^{[\rm ac]}_{\mu\nu}+{h}^{[\rm ac]}_{\mu\nu}\,,\quad\mbox{with}\quad|{h}^{[\rm ac]}_{\mu\nu}|<<1\,.$ (5.1)
@@
Similarly to General Relativity, the branch-cut gravity equations relate the spacetime geometry, encoded in the metric $g^{[\rm ac]}$, to matter described by the energy-momentum tensor $T^{[\rm ac]}_{\mu\nu}$ analytically continued to the complex plane. The Ricci tensor $R^{[\rm ac]}_{\mu\nu}$ and Ricci scalar $R^{[\rm ac]}$ for the linearized theory are computed following the usual scheme, starting from the analytically continued Christoffel symbol which will lead to an analytically continued Riemann curvature tensor. From equation (5.1), following similar mathematical procedures of the standard theory, we obtain for the analytically continued linearized Christoffel symbol
$\Gamma^{{[\rm ac]}\rho}_{\mu\nu}$ $=$ $\frac{1}{2}g^{{[\rm ac]}\rho\sigma}\Bigl{[}\partial_{\mu}g^{[\rm ac]}_{\nu\sigma}+\partial_{\nu}g^{[\rm ac]}_{\mu\sigma}-\partial_{\sigma}g^{[\rm ac]}_{\mu\nu}\Bigr{]}$ (5.3)
@@
$R^{{[\rm ac]}\mu}_{\nu\sigma\rho}$ $=$ $\partial_{\sigma}\Gamma^{{[\rm ac]}\mu}_{\nu\rho}-\partial_{\rho}\Gamma^{{[\rm ac]}\mu}_{\nu\sigma}+\Gamma^{{[\rm ac]}\mu}_{\sigma\lambda}\Gamma^{{[\rm ac]}\lambda}_{\nu\rho}+\Gamma^{{[\rm ac]}\mu}_{\rho\lambda}\Gamma^{{[\rm ac]}\lambda}_{\nu\sigma}\,,$ (5.4)
@@
$R^{[\rm ac]}_{\mu\nu}$ $=$ $R^{{[\rm ac]}\rho}_{\mu\rho\nu}$ (5.5)
$=$ $\frac{1}{2}\Bigl{[}\partial_{\rho}\partial_{\mu}h^{{[\rm ac]}\rho}_{\nu}+\partial_{\rho}\partial_{\nu}h^{{[\rm ac]}\rho}_{\mu}-\partial_{\mu}\partial_{\nu}h^{[\rm ac]}-\Box h_{\mu\nu}\Bigr{]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$
Similarly, the Ricci scalar is
$R^{[\rm ac]}=g^{{[\rm ac]}\mu\nu}R^{[\rm ac]}_{\mu\nu}=\partial_{\mu}\partial_{\nu}h^{{[\rm ac]}\mu\nu}-\Box h^{[\rm ac]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (5.6)
@@
$G^{[\rm ac]}_{\mu\nu}=-\frac{1}{2}\Bigl{[}\Box h^{[\rm ac]}_{\mu\nu}+\zeta^{[\rm ac]}_{\mu\nu}\partial^{\rho}\partial^{\sigma}h^{[\rm ac]}_{\rho\sigma}-\zeta^{[\rm ac]}_{\mu\nu}\Box h^{[\rm ac]}-\partial^{\rho}\partial_{\nu}h^{[\rm ac]}_{\mu\rho}-\partial_{\rho}\partial_{\mu}h^{{[\rm ac]}\rho}_{\nu}+\partial_{\nu}\partial_{\mu}h\Bigr{]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (5.7)
@@
$G^{[\rm ac]}_{\mu\nu}=-\frac{1}{2}\Bigl{[}\Box\bar{h}^{[\rm ac]}_{\mu\nu}+\zeta^{[\rm ac]}_{\mu\nu}\partial^{\rho}\partial^{\sigma}\bar{h}^{[\rm ac]}_{\rho\sigma}-\partial^{\rho}\partial_{\nu}\bar{h}^{[\rm ac]}_{\mu\rho}-\partial^{\rho}\partial_{\mu}\bar{h}^{{[\rm ac]}}_{\nu\rho}\Bigr{]}+{\cal O}({h^{[\rm ac]}}^{2})\,.$ (5.8)
5.2 Stochastic gravitational wave background